OpenStudy (anonymous):

write the equation of a line in point slope form that has a slope of -1 and passes through the point ( -2,5)

OpenStudy (anonymous):

y=mx+b you have "m" (slope) then plug in the x,y and find "b"

OpenStudy (anonymous):

M -1 -5=1(-2)+B

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

7=1+B

OpenStudy (anonymous):

no -5= -1(-2)+b=2+b b= -5-2= -7